OpenStudy (anonymous):

Can someone explain how to differentiate this? i know the answer i just don't understand the process. Find the derivative of the function, f(x)=tan^-1sqrt(3x+1)

OpenStudy (anonymous):

so is st just this particular rpoblem or inverse trig functions in general?

OpenStudy (anonymous):

inverse trig functions in general

OpenStudy (anonymous):

so let's say you have \(y=\sin^{-1}x\). first let's rewrite so we have something we know how to differentiate...\[y=\sin^{-1}x\Rightarrow \sin y = x\] now, differntiate implicitly

OpenStudy (anonymous):

ooh i see, 1/cos is sec. but why does it equal \[\sec(\sin ^{-1}x)\]

OpenStudy (anonymous):

\[(\cos y)\frac{ dy }{ dx }=1\Rightarrow\frac{ dy }{ dx }=\frac{ 1 }{ \cos y }\]

OpenStudy (anonymous):

sorry, you don't want to go that way. (but \(\frac{1}{\cos x}=\sec x\) anyway, we know \(\sin^2 x+ \cos^2 x = 1\Rightarrow \cos x = \sqrt{1-\sin^2 x}\)

OpenStudy (anonymous):

ok i understand that

OpenStudy (anonymous):

so we get \[\frac{ dy }{ dx }=\frac{ 1 }{ \cos y }=\frac{ 1 }{ \sqrt{1-\sin^2 y} }=\frac{ 1 }{ \sqrt{1-\sin^2 (\sin^{-1}x)} }=\frac{ 1 }{ \sqrt{1-x^2} }\]

OpenStudy (anonymous):

remember, \(\sin^2 x = (\sin x)(\sin x) \text{ and } \sin(\sin^{-1} x)=\sin^{-1}(\sin x) = x\)

OpenStudy (anonymous):

i'm trying but i don't understand where \[\sin ^{-1}x\] comes from

OpenStudy (anonymous):

that's y. remember at the beginning that \(y=\sin^{-1}x\)?

OpenStudy (anonymous):

ooh okay got it

OpenStudy (anonymous):

you do the same for all of the inverse trig functions and then use one of the pythagorean identities to get back x, like we did here.

OpenStudy (anonymous):

okay that makes sense, But how does that translate into my original question when there is no y?

OpenStudy (anonymous):

http://www.math.brown.edu/utra/trigderivs.html here is a site that should also help explain the derivatives of inverse trig functions. now, in your question, you have \[f(x)=\tan^{-1}\sqrt{3x+1}\]f(x) is y. anyways, from deriving it yourself or a table of derivatives, \[\frac{ d }{ dx }\tan^{-1}x = \frac{ 1 }{ 1+x^2 }\] you will have to use the chain rule because of the function inside.

OpenStudy (anonymous):

remember, y = f(x)

OpenStudy (anonymous):

yeah i got that i understand up until after \[\frac{ 1 }{ \sqrt{3x+1} } \times {3x+1}\]

OpenStudy (anonymous):

i also use chegg and i have this but i dont understand how they got to the second step and beyond

OpenStudy (anonymous):

\[\frac{ d }{ dx }\tan^{-1}\left( \sqrt{3x+1} \right)=f'\left( x \right)=\frac{ 1 }{1+\left( \sqrt{3x+1} \right)^2 }\cdot \frac{ 1 }{ 2\sqrt{3x+1} }\cdot 3\]

OpenStudy (anonymous):

so\[f'\left( x \right)=\frac{ 1 }{ 1+3x+1 }\cdot \frac{ 1 }{ 2\sqrt{3x+1} }\cdot 3=\frac{ 3 }{ 2\left( 3x+2 \right)\sqrt{3x+1} }\]

OpenStudy (anonymous):

ooh alright, I understand it now. Thank you.

OpenStudy (anonymous):

you're welcome

OpenStudy (anonymous):

if you're good, you should close the question